OpenStudy (anonymous):

The table below compares the radioactive decay rates of two materials. Material Original mass of material (in grams) 1 100 2 200 Mass of material after 11,460 years (in grams) 1 25 2 25 Based on the table, which of these conclusions is most likely correct? A. The half-life of Material 2 is double the half life of Material 1. B. The half-life of Material 1 and Material 2 are equal. C. The half-life of Material 1 is 1,910 years more than the half-life of Material 2. D. The half-life of Material 2 is 1,910 years more than the half-life of Material 1.

OpenStudy (anonymous):

C. is the answer

OpenStudy (anonymous):

can u explain pls

OpenStudy (ash2326):

Half life is the time taken for a radioactive material to get reduced to half of the initial amount. Material 1 initial=100 g , final=25 grams so it has got reduced by a factor of 4 so the period 11460 is actually two half lives of this one so half life of material 1 = 11460/2=5730 years doing the same for material 2 initial=200 g, final=25 grams it's got reduced by a factor of 8, this is equivalent to a period of 3 half lives so half life of material 2= 11460/3=3820 years @vivsss I think you can find the answer now, can't you?